OpenStudy (anonymous):

x − 0.3x = 1,351 how do i solve this?

OpenStudy (anonymous):

Factor out the \(x\): \[x-0.3x=1,351\]\[1x-0.3x=1,351\]\[(1-0.3)x=1,351\]Can you solve this?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

What step confused you?

OpenStudy (anonymous):

Since there is always a 1 in front of the x, add the like terms. (1+.3)x=1351 1.3x=1351 Divide 1.3x and 1351 by 1.3. 1.3x divided by 1.3=1351 divided by 1.3 X=... You can figure out what x is by yourself. hope this helps!!

OpenStudy (anonymous):

I think i just confused you more, i think @wio can explain though

OpenStudy (anonymous):

it just confused me the way you wrote it.... and the last step. the (1-0.3)x=1351 part

OpenStudy (anonymous):

Well can you solve it either way?

OpenStudy (anonymous):

I just dont understand how to answer it

OpenStudy (anonymous):

Find out what x equals.

OpenStudy (anonymous):

Find out what 1-0.3 is, then divide both sides by it.

OpenStudy (anonymous):

0.7

OpenStudy (anonymous):

so its 1930?

OpenStudy (anonymous):

Correct. http://www.wolframalpha.com/input/?i=x+%E2%88%92+0.3x+%3D+1%2C351

OpenStudy (anonymous):

Thank You!!! I get it now lol i knew that ..... brain fart

OpenStudy (anonymous):

It was just distributive property.... x(1-.3) = 1x-.3x But backwards!

OpenStudy (anonymous):

I havent done that kind of problem sense like last year in 7th grade lol i forget things